OpenStudy (anonymous):

Logx base 2 + logx base 3 = 5

OpenStudy (campbell_st):

use change of base \[\log _{2} x = \log _{e}x/\log _{e}2\] \[\log _{3} x = \log _{e} x/\log _{e}3\] so the problem is now \[\log _{e} x/\log _{e} 2 + \log _{e}x/\log _{e}3 = 5\] \[\log _{e} x(1/ \log _{e}2 +1/\log _{e}3) = 5\] \[\log _{e} x = 5/(1/\log _{e}2 + 1/\log _{e}3)\] get a common denominator for the fraction \[\log _{x} = 5/(\log _{e} 3 + \log _{e}2)/(\log _{e}2timeslog _{e}3))\] dividing by a fraction, invert and multiply \[\log _{e} x = 5\times(\log _{e}2 \times \log _{e}3)/(\log _{e}2 + \log _{e}3)\] raise every term to the power of e \[x = e^{(5\ln2\ln3)/(\ln2 + \ln3)}\]

OpenStudy (campbell_st):

this is a tough question becuase of the amount of manipulation needed